# Adaptive Cyclic Diffusion for Inference Scaling

### v2

@@ 3.1 Cyclic Diffusion Search
Noising. The replicated particles are then sent back to an earlier diffusion step $t^{\prime}>0$—referred to as the go-back timestep—via the noising process of the diffusion model: $q({\bf x}_{t^{\prime}}|{\bf x}_{0})$. This operation reintroduces stochasticity and enables the particle to re-enter the denoising trajectory from a different point in the diffusion space. The full cycle—denoising, selection-copy, and noising—is then repeated iteratively until a predefined stopping criterion is met.
@@ 5.7 Do we need to care about go-back temperature level?
Figure 7: (a-b) Per-instance analysis of inference-time scaling behavior across different go-back temperatures. (a) Results for Sudoku instances with 14 provided entities. (b) Results for instances with 44 provided entities. (c) Performance comparison with vs. without adaptive terminal condition in Pixel Maze Size 15. x-axis is the average inference time per sample. (d) Different thinking assignment per hardness in Sudoku (provided number from 17 to 34).
@@ B.5 MoG Result Detail
After leveraging the large go-back to traverse between regions, switching to small go-back becomes advantageous for shifting from exploration to exploitation. This transition enables fine-grained goal-oriented refinement, as demonstrated in Figure 10. All instances begin from an initial incorrect prediction, requiring additional correction during inference. GB=21 fails to escape the initial cluster of modes, whereas GB=81 enables transitions across regions but refines inefficiently. Our method combines coarse transitions to explore distant regions and fine adjustments for local refinement, enabling faster and more effective inference time adjustment.
@@ B.6 OGBench Point Maze Experiment Detail
All methods were evaluated under the same setup, using 32 particles to ensure a fair comparison. For SoP, we used $M=4$ and $K=8$ in Maze Giant, and $M=1$ and $K=32$ in Maze Large. BS used $M=8$ and $K=4$ in both mazes, with a lookahead estimator [16] to have a better predicted $\mathbf{\hat{x}}_{0}(\mathbf{x}_{t})$ with value guidance. SMC [21] was implemented with POTENTIAL TYPE = "sum", $\lambda=0.1$ and used $N=32$ particles in both mazes. ABCD was configured with $N=32$, $K=2$ and $J=16$ in Maze Giant and $N=32$, $K=1$ and $J=32$ in Maze Large. The adaptive terminal condition was set once more than 90% of top-$K$ particles consistently originated from the zero noise level over $\kappa$ consecutive steps. We used $\kappa=30$ in Maze Giant and $\kappa=5$ in Maze Large. All methods used a base diffusion model trained with 256 denoising steps. ABCD employed jumpy denoising with a jump length of 10.
@@ B.7.1 OGBench Point Maze visualization
Figure 11: Visualization of the generated trajectories through iteration with ABCD. The figure highlighted with a red border indicates the selected plan for executing in the environment.
As shown in Figure 11, the initial trajectories proposed by jumpy denoising effectively capture a rough path from the start to the goal, yet contain noticeable invalid segments. Across subsequent iterations, Cyclic Diffusion Search progressively refines these trajectories by focusing on problematic regions and improves them toward viable plans. The highlighted plan selected for execution (red border) showcases ABCD’s capability to preserve the nearly-complete solution by leveraging Automatic Exploration-Exploitation Balancing mechanism.
@@ B.8 Sudoku Experiment Detail
To evaluate inference-time scaling strategies, each baseline was scaled along its primary computational axis while preserving its core algorithmic structure. Base Diffusion employed a single particle ($N=1$). BoN increased inference cost by evaluating models with $N=40$, $480$, and $960$ particles. BS and SoP were implemented with $M=4$ and $K=10$. In BS, inference-time compute was scaled by increasing the branching-and-selection frequency $p$. SoP additionally scaled computation by reducing the step sizes for forward and backward transitions, $(\Delta b,\Delta f)$. SMC used $N=40$ particles and scaled inference-time compute by adjusting the resampling frequency $p$ across timesteps. Following the FKD framework [21], we set the POTENTIAL TYPE to “max” and $\lambda=100$ to enforce an optimization-style behavior. ABCD (ours) was configured with $N=10$, $K=10$, and $J=5$, corresponding to four non-zero go-back noise levels (10, 20, 30, 40) and one zero level. The adaptive terminal condition terminated inference when a specified percentage of top-$N$ particles consistently originated from the zero noise level for $\kappa$ consecutive steps. We fixed max_iter to 100 and scaled ABCD’s inference time by varying both the persistence parameter $\kappa\in{1,2,3,5,10,15,20}$ and percentage thresholds $\in{0.6,0.8,1.0}$. For fast denoising, ABCD employed five total inference steps, each performing a 10-step jump in the diffusion space. Please refer to Table 5 for detailed results.
@@ B.9.1 Sudoku Visualization showing that proper go-back is necessary
To assess inference-time scaling strategies, each baseline was scaled along its primary computational axis while preserving its core algorithmic behavior. Base Diffusion employed a single particle ($N=1$). BoN increased inference cost by using a larger number of particles $N$. BS and SoP were implemented with $M=4$ and $K=10$. In BS, inference-time compute was scaled by increasing the branching and selection frequency $p$. In SoP, additional scaling was achieved by reducing the step sizes for forward and backward transitions, $(\Delta b,\Delta f)$. SMC used $N=40$ particles and scaled inference-time compute by varying the resampling frequency $p$ over steps. Following FKD [21], we used POTENTIAL TYPE = “max” and set $\lambda=100$ to emphasize optimization behavior. ABCD (ours) was configured with $N=10$, $K=10$, and $J=5$, corresponding to four non-zero go-back noise levels (10, 20, 30, 40) and a zero level. The adaptive terminal condition terminated inference once a specified proportion of top-$N$ particles consistently originated from the zero noise level for $\kappa$ consecutive iterations. We scaled ABCD’s inference time by varying $\kappa\in{1,2,3}$ and percentage thresholds $\in{0.1,0.2,0.3,0.6,1.0}$. To ensure computational efficiency, ABCD employed ten denoising steps, each executing a 5-step jump in the diffusion process. Please refer to Table 6 for detailed results.
@@ B.12 Molecular 3D structure Prediction task Experiment Detail
To compare inference-time scaling strategies, each baseline was scaled along its primary computational axis while preserving its core algorithmic behavior. Base Diffusion employed a single particle ($N=1$). BoN was evaluated with using more $N$ to increase inference cost. BS and SoP were implemented with $M=4$ and $K=10$, and inference-time compute was scaled by increasing the branching and selection frequency. Specifically, for BS, we scale inference-time compute by adjusting the branching-and-selection frequency $p$. SoP additionally scaled computation by reducing the step sizes for back-and-forth moves $\Delta f$ and $\Delta b$. SMC used $N=40$ particles and increased inference-time compute by adjusting the resampling frequency $p$ over different steps. Following the FKD [21], we use POTENTIAL TYPE = "max" and $\lambda=100$ to make it as optimization. ABCD (ours) was configured with $N=10$, $K=10$, and $J=5$, corresponding to four non-zero go-back noise levels (50, 100, 200, 400) and a zero level. The adaptive terminal condition halted inference once a specified fraction of top-$N$ particles repeatedly originated from the zero noise level for $\kappa$ consecutive steps. We scaled ABCD’s inference time by varying both $\kappa\in{1,2}$ and percentage thresholds $\in{0.4,0.5}$. To ensure computational efficiency, ABCD employed 100 denoising steps, each executing a 10-step jump in the diffusion process.

### v3

# Adaptive Inference-Time Scaling via Cyclic Diffusion Search

@@ 3.1 Cyclic Diffusion Search
Noising. The replicated particles are then sent back to an earlier diffusion step $0\leq t^{\prime}\leq T$—referred to as the go-back timestep—via the noising process of the diffusion model: $q({\bf x}_{t^{\prime}}|{\bf x}_{0})$. This operation reintroduces stochasticity and enables the particle to re-enter the denoising trajectory from a different point in the diffusion space. The full cycle—denoising, selection-copy, and noising—is then repeated iteratively until a predefined stopping criterion is met.
@@ 5.7 Do we need to care about go-back temperature level?
Figure 7: (a-b) Per-instance analysis of inference-time scaling behavior across different go-back temperatures. (a) Results for Sudoku instances with 14 provided entities. (b) Results for instances with 44 provided entities. (c) Performance comparison with vs. without adaptive terminal condition in Pixel Maze Size 15. x-axis is the average inference time per sample. (d) Different thinking assignment per hardness in Sudoku (provided number from 17 to 33).
@@ B.5 MoG Result Detail
After leveraging the large go-back to traverse between regions, switching to small go-back becomes advantageous for shifting from exploration to exploitation. This transition enables fine-grained goal-oriented refinement, as demonstrated in Figure 10. All instances begin from an initial incorrect prediction, requiring additional correction during inference. go-back $=20$ fails to escape the initial cluster of modes, whereas go-back $=80$ enables transitions across regions but refines inefficiently. Our method combines coarse transitions to explore distant regions and fine adjustments for local refinement, enabling faster and more effective inference time adjustment.
@@ B.6 OGBench Point Maze Experiment Detail
All methods were evaluated under the same setup, using 32 particles to ensure a fair comparison. For SoP, we used $M=4$ and $K=8$ in Maze Giant, and $M=1$ and $K=32$ in Maze Large. BS used $M=8$ and $K=4$ in both mazes, with a lookahead estimator [16] to have a better predicted $\mathbf{\hat{x}}_{0}(\mathbf{x}_{t})$ with value guidance. SMC [22] was implemented with POTENTIAL TYPE = “sum”, $\lambda=0.1$ and used $N=32$ particles in both mazes. ABCD was configured with $N=32$, $K=2$ and $J=16$ in Maze Giant and $N=32$, $K=1$ and $J=32$ in Maze Large. The adaptive terminal condition was set once more than 90% of top-$K$ particles consistently originated from the zero noise level over $\kappa$ consecutive steps. We used $\kappa=30$ in Maze Giant and $\kappa=5$ in Maze Large. All methods used a base diffusion model trained with 256 denoising steps. ABCD employed jumpy denoising with a jump length of 10.
@@ B.7.1 OGBench Point Maze visualization
Figure 11: Visualization of the generated trajectories through cycle with ABCD. The figure highlighted with a red border indicates the selected plan for executing in the environment.
As shown in Figure 11, the initial trajectories proposed by jumpy denoising effectively capture a rough path from the start to the goal, yet contain noticeable invalid segments. Across subsequent cycles, Cyclic Diffusion Search progressively refines these trajectories by focusing on problematic regions and improves them toward viable plans. The highlighted plan selected for execution (red border) showcases ABCD’s capability to preserve the nearly-complete solution by leveraging Automatic Exploration-Exploitation Balancing mechanism.
@@ B.8 Sudoku Experiment Detail
To evaluate inference-time scaling strategies, each baseline was scaled along its primary computational axis while preserving its core algorithmic structure. Base Diffusion employed a single particle ($N=1$). BoN increased inference cost by evaluating models with $N=40$, $480$, and $960$ particles. BS and SoP were implemented with $M=4$ and $K=10$. In BS, inference-time compute was scaled by increasing the branching-and-selection frequency $p$. SoP additionally scaled computation by reducing the step sizes for forward and backward transitions, $(\Delta b,\Delta f)$. SMC used $N=40$ particles and scaled inference-time compute by adjusting the resampling frequency $p$ across timesteps. Following the FKD framework [22], we set the POTENTIAL TYPE to “max” and $\lambda=100$ to enforce an optimization-style behavior. ABCD (ours) was configured with $N=10$, $K=10$, and $J=5$, corresponding to four non-zero go-back noise levels (10, 20, 30, 40) and one zero level. The adaptive terminal condition terminated inference when a specified percentage of top-$N$ particles consistently originated from the zero noise level for $\kappa$ consecutive steps. We fixed max_iter to 100 and scaled ABCD’s inference time by varying both the persistence parameter $\kappa\in\{1,2,3,5,10,15,20\}$ and percentage thresholds $\in\{0.6,0.8,1.0\}$. For fast denoising, ABCD employed five total inference steps, each performing a 10-step jump in the diffusion space. Please refer to Table 5 for detailed results.
@@ B.9.1 Sudoku Visualization showing that proper go-back is necessary
To assess inference-time scaling strategies, each baseline was scaled along its primary computational axis while preserving its core algorithmic behavior. Base Diffusion employed a single particle ($N=1$). BoN increased inference cost by using a larger number of particles $N$. BS and SoP were implemented with $M=4$ and $K=10$. In BS, inference-time compute was scaled by increasing the branching and selection frequency $p$. In SoP, additional scaling was achieved by reducing the step sizes for forward and backward transitions, $(\Delta b,\Delta f)$. SMC used $N=40$ particles and scaled inference-time compute by varying the resampling frequency $p$ over steps. Following FKD [22], we used POTENTIAL TYPE = “max” and set $\lambda=100$ to emphasize optimization behavior. ABCD (ours) was configured with $N=10$, $K=10$, and $J=5$, corresponding to four non-zero go-back noise levels (10, 20, 30, 40) and a zero level. The adaptive terminal condition terminated inference once a specified proportion of top-$N$ particles consistently originated from the zero noise level for $\kappa$ consecutive iterations. We scaled ABCD’s inference time by varying $\kappa\in\{1,2,3\}$ and percentage thresholds $\in\{0.1,0.2,0.3,0.6,1.0\}$. To ensure computational efficiency, ABCD employed ten denoising steps, each executing a 5-step jump in the diffusion process. Please refer to Table 6 for detailed results.
@@ B.12 Molecular 3D structure Prediction task Experiment Detail
To compare inference-time scaling strategies, each baseline was scaled along its primary computational axis while preserving its core algorithmic behavior. Base Diffusion employed a single particle ($N=1$). BoN was evaluated with using more $N$ to increase inference cost. BS and SoP were implemented with $M=4$ and $K=10$, and inference-time compute was scaled by increasing the branching and selection frequency. Specifically, for BS, we scale inference-time compute by adjusting the branching-and-selection frequency $p$. SoP additionally scaled computation by reducing the step sizes for back-and-forth moves $\Delta f$ and $\Delta b$. SMC used $N=40$ particles and increased inference-time compute by adjusting the resampling frequency $p$ over different steps. Following the FKD [22], we use POTENTIAL TYPE = "max" and $\lambda=100$ to make it as optimization. ABCD (ours) was configured with $N=10$, $K=10$, and $J=5$, corresponding to four non-zero go-back noise levels (50, 100, 200, 400) and a zero level. The adaptive terminal condition halted inference once a specified fraction of top-$N$ particles repeatedly originated from the zero noise level for $\kappa$ consecutive steps. We scaled ABCD’s inference time by varying both $\kappa\in\{1,2\}$ and percentage thresholds $\in\{0.4,0.5\}$. To ensure computational efficiency, ABCD employed 100 denoising steps, each executing a 10-step jump in the diffusion process.
@@ Appendix D Prediction visualization
Appendix E Termination guarantee of the Adaptive Thinking Time
In this section, we present a detailed proof establishing the termination guarantee of the Adaptive Thinking Time algorithm introduced in Section 3.3.
E.1 Formal statement
Let
$T=\{t_{1},\dots,t_{m}\}$
be the “temperature pool” with lowest temperature $t_{1}=0$. At cycle $c$, let $D_{k}(c)$ be the multiset of origin-temperatures of the top-$k$ particles after denoising. Define the stopping rule: terminate at the first cycle $N$ such that
$\forall\,i\in\{N-\kappa+1,\,\dots,\,N-1,N\},\quad D_{k}(i)=\{0,0,\dots,0\},$
i.e. all top-$k$ particles originate from $t_{1}=0$ for $\kappa$ consecutive cycles.
Assumptions
  1. 1.
    Bounded reward. There is a reward function $r:\mathbb{R}^{d}\to\mathbb{R}$ with finite supremum $r^{*}<\infty$.
  2. 2.
    Monotonic selection. Let
    $r_{k}(c)\;=\;\min\{\text{rewards of top-}k\text{ at cycle }c\}.$
    Then $r_{k}(c)$ is non-decreasing in $c$.
  3. 3.
    Attainment. There exists at least one state $x^{*}$ such that $r(x^{*})=r^{*}$.
Theorem 1.
Under these assumptions, ABCD’s Adaptive Thinking Time criterion triggers termination in finite time.
The proof of this theorem is built upon several key lemmas, which we will now introduce and prove.
E.2 Uniform Hit-Chance
Lemma E.1.
Let $x^{*}$ be any maximizer of the reward, $r(x^{*})=r^{*}.$ Under the ABCD dynamics (Algorithm 1), there is a constant $p>0$ such that for each cycle $c$,
$\Pr\bigl{(}E_{c}\mid\mathcal{F}_{c-1}\bigr{)}\;\geq\;p,$
where $E_{c}=\text{``$x^{*}$ appears among the selected top-}K\text{ at cycle $c$\textquotedblright}$ ” and $\mathcal{F}_{c-1}$ is the $\sigma$-algebra of all randomness up to (but not including) cycle $c$.
Proof.
  1. 1.
    At the start of cycle $c$ we have $K$ anchor particles at $t=0$. We replicate each $J$ times and send each of those $KJ$ copies through the forward noising process $q(x_{t^{\prime}}\mid x_{0})$ at every $t^{\prime}\in T$. Exactly $K$ of those replicas go to the highest temperature $t_{m}$.
  2. 2.
    Conditioned on $\mathcal{F}_{c-1}$, the noising and denoising of these $KJ$ particles is independent of the past and identically distributed across cycles. In particular, the $K$ samples at $t_{m}$ are fresh each time. Let
    $q\;=\;\Pr\bigl{[}\text{(noising at }t_{m}\text{) then (denoise) yields }x^{*}\bigr{]}.$
    By full support of the forward kernel and nonzero recovery probability of the denoiser, $q>0$.
  3. 3.
    Those $K$ draws at $t_{m}$ are independent, so the chance none equals $x^{*}$ is $(1-q)^{K}$. Hence
[…]
  4. 4.
    If any replicate equals $x^{*}$, its reward is $r^{*}$, guaranteeing $x^{*}$ is in the top-$K$. Thus for every history,
    $\Pr\bigl{(}E_{c}\mid\mathcal{F}_{c-1}\bigr{)}\;=\;\Pr\bigl{(}\text{some }t_{m}\text{--replicate}=x^{*}\mid\mathcal{F}_{c-1}\bigr{)}\;\geq\;p.$
This completes the proof of the uniform hit-chance lemma. ∎
E.3 Optimality
Lemma E.2.
Under the assumptions
  1. 1.
    Monotonicity and Boundedness: $r_{k}(c)$ is non-decreasing and $r_{k}(c)\leq r^{*}<\infty$, so $r_{k}(c)\to r_{\infty}\leq r^{*}.$
  2. 2.
    Supremum Attained: $\exists\,x^{*}$ with $r(x^{*})=r^{*}$.
  3. 3.
    Uniform Hit-Chance: at each cycle $c$, $\Pr(E_{c}\mid\text{past})\geq p>0.$
Then
[…]
Proof.
  1. 1.
    Define the “hit” event $E_{c}$ as above. By assumption 3, $\sum_{c=1}^{\infty}\Pr(E_{c}\mid\text{past})\geq\sum_{c}p=\infty.$
  2. 2.
[…]
  3. 3.
    Let $C$ be the first cycle where $E_{C}$ occurs. Then
    $r_{k}(C)\geq r(x^{*})=r^{*},\quad r_{k}(C)\leq r^{*},\quad\implies\quad r_{k}(C)=r^{*}.$
  4. 4.
    Monotonicity then implies $r_{k}(c)=r^{*}$ for all $c\geq C$, so $r_{\infty}=r^{*}$.
Hence $\Pr[r_{\infty}=r^{*}]=1.$
E.4 Finite-Time Hitting
Lemma E.3.
Under the Uniform Hit-Chance lemma, define
$T\;=\;\inf\bigl{\{}\,c\geq 1\mid x^{*}\text{ appears among the top-}K\text{ atcycle }c\bigr{\}}.$
Then
$\Pr\bigl{(}T<\infty\bigr{)}=1,\quad\Pr\bigl{(}T>n\bigr{)}\leq(1-p)^{n}\quad\forall\,n\in\mathbb{N},$
where $p>0$ is the per-cycle success lower bound.
Proof.
  1. 1.
    Failure-to-hit bound.
    Let
    $F_{n}\;=\;\bigcap_{c=1}^{n}E_{c}^{c},\quad E_{c}=\{\text{cycle }c\text{ hits }x^{*}\}.$
    By the Uniform Hit-Chance lemma,
    $\Pr\bigl{(}E_{c}\mid\mathcal{F}_{c-1}\bigr{)}\geq p\quad\Longrightarrow\quad\Pr\bigl{(}E_{c}^{c}\mid\mathcal{F}_{c-1}\bigr{)}\leq 1-p.$
    Hence, by the tower property,
[…]
    and by induction
    $\Pr(F_{n})\leq(1-p)^{n}.$
  2. 2.
    Hitting in finite time w.p.1.
    Since $\{T>n\}=F_{n}$, we get
[…]
    so $\Pr(T=\infty)=0$ and hence $\Pr(T<\infty)=1.$
E.5 Proof of Main Theorem
We now assemble the results from the lemmas to prove Theorem 1.
  1. 1.
    Monotone convergence of $r_{k}(c)$. The sequence $r_{k}(0),r_{k}(1),r_{k}(2),\dots$ is non-decreasing and bounded above by $r^{*}$, hence it converges to some limit $r_{\infty}\leq r^{*}$.
  2. 2.
    Limit must equal $r^{*}$. According to Lemma E.2, $r_{\infty}=r^{*}$.
  3. 3.
    Optimal is reached in finite time. By Lemma E.3, $r_{k}(c)$ hits $r^{*}$ at some finite cycle $C$.
  4. 4.
    Once optimal is reached, only $t_{1}$-origins survive. At cycle $C$ where $r_{k}(C)=r^{*}$, any particle noised at temperatures $t>0$ and then denoised back cannot exceed $r^{*}$. Thus the only way to preserve the top-$k$ set of reward-$r^{*}$ particles is via replicas from $t_{1}=0$, which leave them unchanged. Consequently, for all $c\geq C$, $D_{k}(c)=\{0,0,\dots,0\}$.
  5. 5.
    $\kappa$-persistence triggers termination. Therefore, once cycle $C$ is reached, the condition “all top-$k$ from temperature 0 for $\kappa$ consecutive cycles” is satisfied by cycle $C+\kappa-1$, and ABCD terminates.